OpenStudy (anonymous):

Solve the equation using factoring. x^2 + 7x + 12 = 0 {2, 6} {-2, -6} {3, 4} {-3, -4}

OpenStudy (anonymous):

(x+3)(x+4)=0 x=-3,-4

OpenStudy (anonymous):

x2+4x+3x+12=0 x(x+4)+3(x+4()=0 x+3)(x+4)=0 x=-3 x=-4

OpenStudy (anonymous):

x = -3 or x=-4